Systems Staff Meeting Notes 6/21/99

Systems Staff Meeting Notes
June 21, 1999

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Suzy Shaw, Ying Tang
Announcements
  1. Preservation/Digital Library shuffle - Erich Kesse and the Digital Library will be taking over half of LAC, including room 419 (currently used to store equipment destined for survey). The proposed bookstore will move from the 2nd to the 1st floor, occupying a small notch of the Preservation Department. Surveyed equipment will be housed in the area in Preservation where they currently store newspapers. This is the area where Gus currently works. It will be under Intellikey. (If the equipment currently housed in 419 is not picked up by Property before the time of the move, Systems will be responsible for moving everything into the new storage area). Preservation's newspaper area will move to the 2nd floor - where the bookstore was originally going to be located.
    The North Tower will house a single server that will handle all of the Digital Library, Preservation & Resource Services. These moves may require some wiring, but we should have enough in the tower.
    As of July 1, Erich Kesse will no longer be the head of Preservation - he will be the head of the Digital Library, reporting to Dale. The Digital Library will be separate from Preservation.
  2. Security - The vault combinations have been changed, only Bill Covey and the networking staff know them. When closing up the vaults, take care to spin the dial multiple times to insure proper closure. The Visitor Log seems to be working. Everyone, even the housekeeping staff, needs to be signed in & out of the office. The door latch on room 10 (front door) has been repaired and seems to be functioning properly now. There has been no word from UPD regarding the theft.
  3. NERDC Account Funding - the master was at zero this morning. Bill will juggle the funding by hand. We will not do the normal funding amount in the morning. If anyone calls about a lack of money, alert Bill. We have cut the EDP funding by a factor of 6. The Dean's Menu is now one of our accounts. SAS in ILL costs us the most. Winston is showing major EDP spending, will check it for accuracy.
  4. Christian will be moving to Tigert, then on to Yon Hall. Housekeeping would like a schedule for cleaning the web suite. Suzy will come up with one.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Servers received, one has been installed.
  2. Cow shipment is due this week.
  3. Switches were received last week. Everyone should eventually be on an individual switch port. Will start in Systems, West to follow.
Liaisons
  1. No date established yet for the training session.
Networking
  1. Mail problem today - related to password changes on Dewey. Winston changed the password back to the old one so that the mail would come up, but we will need to try and successfully change that password today.
Software Projects Status
  1. Suzy attended InterDev training last week. InterDev is a web environment that works with OBDC type databases. It provides good support for creating HTML and scripting languages. It enables teamwork on web pages with auto-management. There were only 5 people in attendance, one was a 14 year-old who had already passed most of the Microsoft tests. InterDev is a software suite - we have one copy in Systems - and is an easier hook for the Microsoft 3-tier software. Mainly NTS, it supports JAVA, VisualBasic, etc. Maybe useful in the future if we have teams for web development. Overall, Suzy found it to be a good experience.
  2. Plan from the CIO - EAGLE - hooks a mainframe CICS application to the web. It will be used by all of the big CICS databases. CCI (Community Computing Initiative) will produce data warehouses for general use on campus. EAGLE can be used as an interface - OBDC would work for us. The plan is 28 pages long & covers most computing on campus.
  3. Winston spent much of the week recovering from the server change - SQL to NT security. He is still dealing with a CrystalReports problem. (He may have to rewrite with QuickReports, except for in Accounting.) Ying's sign-on is still problematic due to the server upgrade - Winston & Microsoft have been back & forth on the phone - he recently FTP'd the master file to Microsoft so they can recreate the problem.
  4. All new Library web pages have been moved to the NT8 server. Rich is planning to activate them & advertise them on Caroline by tomorrow. This should work out nicely, providing ample time for feedback by the end of Summer B and for the necessary tweaking to be in place before the Fall semester. NT8 has been upgraded - it now has 256MB of memory, two Pentium Pro Processors & has been optimized for NT.
  5. Security hole: the .HTR extension used on servers allowed for passwords to be changed via the web interface. We've disabled that option on the web & mail servers, a patch will be available this week & we will install it.
  6. Baldwin is finished, all records have been inputted to FCLA.
DP Coordinator
  1. University of Iowa was able to read our tapes. They've shipped them back to us, upon receipt we will take them to Bevis at NERDC.
  2. Brian & Leo will be working this week (break week).
Calendar
  1. Everett out today.
  2. Rich out June 25-29 (ALA).
  3. Winston out July 1 & 2.
  4. Everett will be out July 2-9.
  5. Bobbie, Ying and Winston out July 12-16.
  6. Rich out July 19-30.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu